A train is traveling at a speed of 26 m/s. An inconsiderate passenger tosses an apple core directly out the window (perpendicular to the side of the train) at a velocity of 3.3 m/s relative to the train. Find the angle that the apple core’s velocity makes with the track as viewed by a wolf sitting nearby.
1
Expert's answer
2019-09-24T09:08:32-0400

"\alpha = \arctan\frac{v_{apple-train}}{v_{train}} = \arctan \frac{3.3 \text{ m/s}}{26.0 \text { m/s}} \approx 7.23 \degree."